# A lattice study of two dimensional SU(2) gauge theories with a single massless Majorana fermion

### v1

@@ I Introduction and summary
The fermion bilinear condensate in the limit of $T\to 0$ is defined as the expectation value of the mass operator in the ground state ($E=E_{0}$) at a fixed $z$ in Dempsey et al. (2024). If the condensate is due to the presence of exact zero modes, it cannot be calculated as an expectation value with respect to a single partition function defined above but we can consider [^2] The need for summing over different twist sectors has been proposed in four dimensional SYM Anber and Poppitz (2025).
@@
can be non-zero and different at finite temperature or as $T\to 0$. The above argument translates to partition functions on an $\ell^{2}$ torus where $\ell$ is the finite extent of the torus and we use the Lagrangian formalism to study the theory. We will refer to the density associated with the quantities in Eq. (7) as topological bilinear condensates since they are present on a finite torus due to the mod($2$) index and the mass operator is the ’t Hooft vertex that absorbs the zero modes. In other words, the $Z_{2}$ chiral symmetry is explicitly broken by the presence of the zero modes. The above argument does not address the issue of the fermion bilinear condensate arising from spontaneous symmetry breaking. Consider
@@
    The spectral density of the eigenvalues of the Majorana fermion indicates the presence of a $\delta$-function at zero eigenvalue in sectors where there are no zero modes. This suggests the need to sum over partition functions to properly reproduce clustering property. Futhermore, there is no evidence for an additional final spectral density at zero eigenvalue. In addition, the ordered eigenvalues given by
@@ III Details of the lattice formalism
For the range of $\ell$ we will need to study the physics of interest to us, we found it optimal to generate gauge fields using just the gauge action in Eq. (45) and treat the fermionic determinant as an observable  [^4] This method is inspired by the results obtained for the fermionic condensate in the Schwinger model using a similar method Narayanan et al. (1995).. Since we wanted to ensure that the set of gauge fields that were generated were independent, each configuration was obtained from a random cold start that was thermalized using the heat-bath algorithm Kennedy and Pendleton (1985) till it was thermalized. In order to ensure a proper covering of all configurations that extremizes the action for $z=1$, we set all $U^{g}_{k}(n_{1},n_{2})=1$ except
@@
The partition functions $Z_{zf}$ with $(zf)=(++)$ and (zf)=(+)(zf)=(+-) corresponds to setting $z=1$ in Eq. (45) and imposing periodic or anti-periodic boundary conditions of fermions and (zf)=(+)(zf)=(-+) and (zf)=()(zf)=(--) corresponds to setting z=1z=-1 in Eq. (45) and imposing periodic or anti-periodic boundary conditions of fermions.
@@ V A mod($2$) topological condensate in finite volume
and we see that it approaches unity as $\ell\to\infty$. The results for $\Sigma_{t,\pm,1}(\ell,\infty)$ are shown in the bottom-right panel of Figure 2. We have only shown the results after extrapolation to $L\to\infty$ to aviod cluttering the plot. The behavior at finite $\ell$ are qualitatively different from the ones in the other panels but the result as $\ell\to\infty$ is numerically consistent with the ones in the other panels. The main contribution to the difference at finite $\ell$ arises from the $\ell$ behavior of Eq. (65).
@@ VI Absence of spontaneous symmetry breaking
The values for $\gamma_{m}(\pm 1,1)$ are consistent with each other. The higher statistics at $(z,J)=(1,1)$ shows that a value of 1-1 is ruled out by $4.5$ standard deviations. A value of 1-1 at $(z,J)=(1,2)$ is ruled out by $3.3$ standard deviations even with an error that is larger by a factor of $2$ compared to $(z,J)=(1,1)$. Finally a value of 1-1 is clearly ruled out for $(z,J)=(1,3),(1,4)$. In addition we also see a decreasing trend in $\gamma_{m}$ as $J$ is increased. We should conclude our analysis of Figure 3 and Figure 4 by pointing that the results for $(z,J)=(0,1)$ shown in the bottom-right panel of both the figures behave in a manner similar to the other five panels and result in values for $\gamma_{m}$ consistent with the other two values for $J=1$.
@@ VII Discussion and conclusions
The results obtained in this paper are counter-intuitive. Gauge field boundary conditions or fermion boundary conditions are usually not expected to affect the final physics results – they are usually used to suppress finite volume effects or control the effects of trivial fermion zero modes. We have shown that it drastically affects the partition function in the continuum limit to the extent that it is not possible to show evidence for a non-zero fermion bilinear condensate by conventional methods. One way out is to sum over boundary conditions of gauge fields noting that it is not the same as summing over topological sectors although the presence of the mod($2$) index suggests an analogy. We have the option of summing over fermion boundary conditions or gauge field boundary conditions. Since it might be more natural to sum over gauge field boundary conditions with fixed fermion boundary conditions, we may consider the standard gauge action with the extra parameter $z$ in Eq. (45) not necessarily set to $\pm 1$. In a two dimensional non-abelian gauge action, the non-trivial constraint on the Wilson loop obtained by winding on both directions of the torus can be localized to the flux at a point on the torus. The translation to a Wilson gauge action is simple – one of the $L^{2}$ plaquettes on the lattice is responsible for the global Wilson loop that winds around both directions of the torus. If we set the coupling of that plaquette to zero ($z=0$ in Eq. (45)), then we have an action with independent plaquettes and a measure on the global Wilson loop that is independent of the gauge coupling. How can changing the contribution of one plaquette to the lattice action matter in the continuum limit? Contrary to the standard actions with $z=\pm 1$, where all gauge field configurations either produce zero modes or not depending on the boundary conditions, action with $z=0$ results in a fraction $f$ of the configurations to have zero modes or not depending on the boundary conditions and this fraction changes from a value away from $\frac{1}{2}$ at small volumes to $\frac{1}{2}$ at large volumes. This difference is seen when one compares the bottom-right panel ($z=0$) of Figure 1 to the other five panels ($z=\pm 1$). With $z=0$, we can compute the topological condensate without mixing boundary conditions. We find that such a condensate is non-zero at all sizes of the continuum torus for both choices of boundary conditions of fermions implying that the symmetry is broken explicitly.

### v3

# Lattice study of two-dimensional SU(2) gauge theories with a single massless Majorana fermion

@@ I Introduction and summary
The fermion bilinear condensate in the limit of $T\to 0$ is defined as the expectation value of the mass operator in the ground state ($E=E_{0}$) at a fixed $z$ in Dempsey et al. (2024). If the condensate is due to the presence of exact zero modes, it cannot be calculated as an expectation value with respect to a single partition function defined above but we can consider [^2] Mixed partition functions were also considered in the context of four dimensional SYM Anber and Poppitz (2023, 2025a, 2025b) to enable a computation of the gluino condensate by mixing twisted sectors.
@@
can be non-zero and different at finite temperature or as $T\to 0$. The above argument translates to partition functions on an $\ell^{2}$ torus where $\ell$ is the finite extent of the torus and we use the Lagrangian formalism to study the theory. We will refer to the density associated with the quantities in Eq. (7) as topological bilinear condensates since they are present on a finite torus due to the mod($2$) index and the mass operator is the ’t Hooft vertex that absorbs the zero modes. In other words, the $Z_{2}$ chiral symmetry is explicitly broken by the presence of the zero modes. We will refer such a condensate as a topological condensate. This is different from a fermion bilinear condensate arising from spontaneous symmetry breaking. Consider
@@
    One could study spontaneous symmetry breaking by introducing a mass term and studying the massless limit. Such an approach requires one to first establish the small mass range and then study the behavior in that range. A better option from the numerical perspective is to study the spectral density which carries the same information. The spectral density of the eigenvalues of the Majorana fermion indicates the presence of a $\delta$-function at zero eigenvalue in sectors where there are no zero modes. This suggests the need to sum over partition functions to properly reproduce clustering property. Furthermore, there is no evidence for an additional final spectral density at zero eigenvalue. In addition, the ordered eigenvalues given by
@@ III Details of the lattice formalism
For the range of $\ell$ we will need to study the physics of interest to us, we found it optimal to generate gauge fields using just the gauge action in Eq. (45) and treat the fermionic determinant as an observable  [^4] This method is inspired by the results obtained for the fermionic condensate in the Schwinger model using a similar method Narayanan et al. (1995).. It is sufficient to use a standard heat-bath algorithm to generate a sequence of configurations. Since we want to approach small values of $\ell$, we need to ensure that the zero modes of the gauge fields are properly sampled. We also need to ensure that the set of gauge fields that were generated were independent. To this end, each configuration was obtained from a random cold start that was thermalized using one heat-bath step followed by one over-relazation step. We used Kennedy and Pendleton (1985) for the heat-bath step and Creutz (1987); Adler (1981, 1988); Neuberger (1987) for the overrelaxation step. We ran the algorithm for enough pairs of heat-bath and overrelaxation steps to ensure that we are well into the thermalized regime. We could have continued taking measurements after thermalization but to ensure that the starting point was well sampled for the zero modes of gauge fields, we restarted after every measurement and ensured thermalization for each measurement. Since the main cost of the computation was in the evaluation of the fermionic eigenvalues, we could afford to add this minimal extra cost. We set all $U^{g}_{k}(n_{1},n_{2})=1$ except
@@
This was the procedure for generating the starting configuration. The partition functions $Z_{zf}$ with $(zf)=(++)$ and (zf)=(+)(zf)=(+-) corresponds to setting $z=1$ in Eq. (45) and imposing periodic or anti-periodic boundary conditions of fermions and (zf)=(+)(zf)=(-+) and (zf)=()(zf)=(--) corresponds to setting z=1z=-1 in Eq. (45) and imposing periodic or anti-periodic boundary conditions of fermions. We also ensured that we are in the correct continuum phase where Polyakov loops in both directions had averages consistent with zero.
@@ V A mod($2$) topological condensate in finite volume
and we see that it approaches unity as $\ell\to\infty$. The results for $\Sigma_{t,\pm,1}(\ell,\infty)$ are shown in the bottom-right panel of Figure 2. We have only shown the results after extrapolation to $L\to\infty$ to avoid cluttering the plot. The behavior at finite $\ell$ are qualitatively different from the ones in the other panels but the result as $\ell\to\infty$ is numerically consistent with the ones in the other panels. The main contribution to the difference at finite $\ell$ arises from the $\ell$ behavior of Eq. (65).
@@ VI Absence of spontaneous symmetry breaking
The values for $\gamma_{m}(\pm 1,1)$ are consistent with each other. The higher statistics at $(z,J)=(1,1)$ shows that a value of $1$ is ruled out by $4.5$ standard deviations. A value of $1$ at $(z,J)=(1,2)$ is ruled out by $3.3$ standard deviations even with an error that is larger by a factor of $2$ compared to $(z,J)=(1,1)$. Finally a value of $1$ is clearly ruled out for $(z,J)=(1,3),(1,4)$. In addition we also see a decreasing trend in $\gamma_{m}$ as $J$ is increased. We should conclude our analysis of Figure 3 and Figure 4 by pointing that the results for $(z,J)=(0,1)$ shown in the bottom-right panel of both the figures behave in a manner similar to the other five panels and result in values for $\gamma_{m}$ consistent with the other two values for $J=1$.
@@ VII Discussion and conclusions
The results obtained in this paper are counter-intuitive. Gauge field boundary conditions or fermion boundary conditions are usually not expected to affect the final physics results – they are usually used to suppress finite volume effects or control the effects of trivial fermion zero modes. We have shown that it drastically affects the partition function in the continuum limit to the extent that it is not possible to show evidence for a non-zero fermion bilinear condensate by fixing the boundary conditions for gauge and fermion fields. One way out is to sum over boundary conditions of gauge fields noting that it is not the same as summing over topological sectors although the presence of the mod($2$) index suggests an analogy. We have the option of summing over fermion boundary conditions or gauge field boundary conditions. Since it might be more natural to sum over gauge field boundary conditions with fixed fermion boundary conditions, we may consider the standard gauge action with the extra parameter $z$ in Eq. (45) not necessarily set to $\pm 1$. In a two dimensional non-abelian gauge action, the non-trivial constraint on the Wilson loop obtained by winding on both directions of the torus can be localized to the flux at a point on the torus. The translation to a Wilson gauge action is simple – one of the $L^{2}$ plaquettes on the lattice is responsible for the global Wilson loop that winds around both directions of the torus. If we set the coupling of that plaquette to zero ($z=0$ in Eq. (45)), then we have an action with independent plaquettes and a measure on the global Wilson loop that is independent of the gauge coupling. How can changing the contribution of one plaquette to the lattice action matter in the continuum limit? Contrary to the standard actions with $z=\pm 1$, where all gauge field configurations either produce zero modes or not depending on the boundary conditions, action with $z=0$ results in a fraction $f$ of the configurations to have zero modes or not depending on the boundary conditions and this fraction changes from a value away from $\frac{1}{2}$ at small volumes to $\frac{1}{2}$ at large volumes. This difference is seen when one compares the bottom-right panel ($z=0$) of Figure 1 to the other five panels ($z=\pm 1$). With $z=0$, we can compute the topological condensate without mixing boundary conditions. We find that such a condensate is non-zero at all sizes of the continuum torus for both choices of boundary conditions of fermions implying that the symmetry is broken explicitly.